OpenStudy (anonymous):

Using quotient rule, how do i set up j(x)= (3x + cosx)/ f(x) to find j'(x)?

OpenStudy (caozeyuan):

\[(\frac{ u }{ v })\prime=\frac{ u \prime v-v \prime u }{ v^{2} }\]

OpenStudy (caozeyuan):

where u and v are both functions of x

OpenStudy (caozeyuan):

so what is u and v in your question?

OpenStudy (anonymous):

j(x)= (3x + cosx) divided by f(x) I'm not sure if I have to separate them, like j(x)= (3x/f(x)) + (cosx/f(x)) and use both product and quotient rule

OpenStudy (caozeyuan):

nonono, just plug the function in to the equation

OpenStudy (caozeyuan):

so your u is 3x + cos x and v is f(x)

OpenStudy (anonymous):

is this right? j'(x)= f(x)(3-sinx) - (3x+cosx)f'(x) divided by [f(x)]^2

OpenStudy (caozeyuan):

well I think this is final answer

OpenStudy (anonymous):

ok thanks!!